OpenStudy (anonymous):

solve for y x=y/(y-1)

OpenStudy (anonymous):

ok i cant just give you the answer so ask a question or something

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

can you assist me in walking through it?

OpenStudy (anonymous):

well let me see

OpenStudy (anonymous):

there are two ways

OpenStudy (anonymous):

I get (y-1)x=y and i dont know what to do after that

OpenStudy (anonymous):

check it out

OpenStudy (anonymous):

multiply them xy-x = y

OpenStudy (anonymous):

did it help?

OpenStudy (anonymous):

its telling me to create an account

OpenStudy (anonymous):

xy-y = x y(x-1) = x y= x/x-1

OpenStudy (anonymous):

reaally i thought if you just go in the link it will tell you something else let me see

OpenStudy (anonymous):

y=x/x-1

OpenStudy (anonymous):

thats what the page was supposed to say

OpenStudy (anonymous):

so the starting equation was y=x/(x-1) and the inverse is the same?

OpenStudy (anonymous):

looks like it

OpenStudy (anonymous):

but the graph would look different.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

another way : x=y/y-1 ... flip them 1/x=(y-1)/y 1/x=y/y - 1/y ..... spread them 1/x = 1- 1/y 1/y=1-1/x 1/y = (x-1)/x then y = x/x-1

OpenStudy (anonymous):

about the graph ... i think so

OpenStudy (anonymous):

like the point (3,1.5) should be (1.5,3)

OpenStudy (anonymous):

thats why im so confused. I am getting the same equation that is supposed to be a different graph

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

this is one equation that we are talking about

OpenStudy (anonymous):

sorry. i wasnt detailed in the original question. i needed to find the inverse of y=x/(x-1)

OpenStudy (anonymous):

i just posted where i was getting stuck

OpenStudy (anonymous):

did you get the inverse?

OpenStudy (anonymous):

I cant figure it out. i was searching online and found one source that said it had no solution. I have no idea what it is

OpenStudy (anonymous):

hmmm i think there is an inverse

OpenStudy (anonymous):

thats what i was thinking as well. I am sure there is one. but its confusing

OpenStudy (anonymous):

can you not just switch the y (s) for x ?

OpenStudy (anonymous):

well i need to solve for y

OpenStudy (anonymous):

then it becomes a circle

OpenStudy (anonymous):

well this is quite the confusing problem then.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

try to sketch the two graphs

OpenStudy (anonymous):